Solve the equation
1) -4[3x + 7 + 3(x + 1)] = -7x - 5
2) 4x/5 -x =5/6-(x-1/4)

Found 2 solutions by rfer, MathTherapy:
Answer by rfer(16322)

1)
-4[3x+7+3(x+1)]=-7x-5
-4[3x+7+3x+3]=-7x-5
-4(6x+10)=-7x-5
-24x-40=-7x-5
-24x+7x=-5+40
-17x=35
x=35/-17
x=2.06
------------------
2)
4x/5-x=5/6-(x-1/4)
4x/5-x=5/6-x+1/4
4x/5-5x/5-5x/5=5/6+1/4
4x/5=10/12+3/12
4x=(13/12)
x=(13/12)/(4/1)
x=13/12*1/4
x=13/48

Answer by MathTherapy(10552)


Solve the equation
1) -4[3x + 7 + 3(x + 1)] = -7x - 5
2) 4x/5 -x =5/6-(x-1/4)